OpenStudy (anonymous):

Solve the porportion, When necessary round to the nearest tenth. The question below. A medal to anyone who helps.

OpenStudy (anonymous):

A. 13 B. 1 C. 7 D. 16

OpenStudy (anonymous):

First cross multiply, do you know how to do that with this problem?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

8 x 4 = ? 2(x + 3) = ?

OpenStudy (anonymous):

The answer is D

OpenStudy (anonymous):

32 = 2x + 6 -6 - 6 26 = 2x 2 2 x = 13

OpenStudy (anonymous):

I meant A. I put the wrong letter.

OpenStudy (anonymous):

13 + 3 + 16. 16/4 =4 and 8/2 = 4